Data for Barry Computer Co. and its industry averages follow. The firm's debt is priced at par, so the market value of its debt equals its book value. Since dollars are in thousands, number of shares are shown in thousands too.

Barry Computer Company:
Balance Sheet as of December 31, 2018 (In Thousands)
Cash $90,720 Accounts payable $196,560
Receivables 589,680 Other current liabilities 151,200
Inventories 302,400 Notes payable to bank 75,600
   Total current assets $982,800    Total current liabilities $423,360
Long-term debt $302,400
Net fixed assets 529,200 Common equity (78,624 shares) 786,240
Total assets $1,512,000 Total liabilities and equity $1,512,000
Barry Computer Company:
Income Statement for Year Ended December 31, 2018 (In Thousands)
Sales $2,100,000
Cost of goods sold
   Materials $903,000
   Labor 546,000
   Heat, light, and power 105,000
   Indirect labor 168,000
   Depreciation 84,000 1,806,000
Gross profit $ 294,000
Selling expenses 126,000
General and administrative expenses $ 42,000
   Earnings before interest and taxes (EBIT) $ 126,000
Interest expense 36,288
   Earnings before taxes (EBT) $ 89,712
Federal and state income taxes (40%) 35,885
Net income $ 53,827
Earnings per share $ 0.68461
Price per share on December 31, 2018 $ 11.00
  1. Calculate the indicated ratios for Barry. Round your answers to two decimal places.
    Ratio Barry              Industry Average
    Current x 2.27x
    Quick x 1.64x
    Days sales outstandinga days 47.89 days
    Inventory turnover x 7.58x
    Total assets turnover x 1.65x
    Profit margin   % 2.40%
    ROA   % 3.96%
    ROE   % 7.61%
    ROIC   % 7.50%
    TIE x 3.50x
    Debt/Total capital   % 33.57%
    M/B   % 4.10%
    P/E   % 18.51%
    EV/EBITDA   % 7.89%

    aCalculation is based on a 365-day year.
  2. Construct the DuPont equation for both Barry and the industry. Round your answers to two decimal places.
    FIRM INDUSTRY
    Profit margin   % 2.40%
    Total assets turnover x 1.65x
    Equity multiplier x x
Answer #1

As per rules I am answering the first 4 subparts of the question

1: Current ratio = Current assets/ Current liabilities

= 982800/423360

=2.32

2: Quick ratio = (Cash + receivables)/ current liabilities

= (90720+589680)/ 423360

=1.61

3: DSO= AR*365/ Sales

= 589680*365/ 2100000

=102.49 days

4: Inventory turnover = COGS/ Average inventory

= 1806000/ 302400

=5.97